# Information That Matters: Exploring Information Needs of People Affected by Algorithmic Decisions

### v1

@@ 4. Results
Questions on the data used by the system emerged in both use cases (43 questions total) and occurred mostly in the second phase of inquiry after reference to the XAI Question Bank (Liao et al., 2020). The question bank already contains questions about data type (What kind of data was the system trained on?) and data source (What is the source of the training data?). As multiple participants asked about data privacy and misuse (11 questions by 8 participants), we add a subcategory on data safety: Are the data safe?
4.1.2. System context: Intention, deployment process, consequences, development, target group, responsibility, perspective, and values.
This category captures questions that aimed to elicit background information on various facets influencing the system’s development and deployment (69 total).
Intention: What is the intention of deploying it? Inquiries about the intention of the system’s deployment emerged in both use cases, with a majority in employment prediction (12 of 18 total). These questions mostly aimed at the intention of deploying the system, which we differentiate regarding i) practical intention, such as data collection, ”How did they come up with the idea of GPS? What’s the intention behind it?” (B4); and ii) structural intention, such as societal benefit, ”What is the benefit for society? What did they expect from this algorithm?” (A1).
Deployment process: How does the deployment process work? Inquiries about the general deployment process emerged mostly in the employment prediction use case (13 of 19 total), including how the system should be made available, if it had already been tested, and how comparable systems were deployed in other countries.
Consequences: What are the consequences after deployment? Inquiries about effects that could follow from the system’s deployment emerged in low counts in both use cases (9 total). Topics of inquiry included, e.g., the effect on the available workforce and care ratio, and often had a focus on risks and challenges:
Won’t I lose labor potential if I only decide based on educational background and so on? If this algorithm is introduced, doesn’t it lead to the loss of workforce? (A1)
Development: How is the system developed? Questions on the system’s technical development emerged mainly in the employment prediction use case (8 of 9 total) and covered, e.g., the contracted company, duration of development, and the collaboration between developers and domain experts. These questions were asked primarily by personnel consultants, who related the information to their own experience with contracted software developers and compared it to known procedures.
Target group: Who is the system’s intended target group? Questions on the system’s target groups emerged mainly in the wristband use case, were low in number (8 of 10 total), and asked mostly about the individual background of the target group. In contrast, the 2 inquiries in the employment prediction use case focused on target groups on a societal level: ”Would the system be applied for every population group? Can we say that every citizen would then be covered?” (A11).
Responsibility: Who is responsible for the system’s deployment? Four questions, two in each use case, addressed who would be responsible for the system’s deployment. This low count is noteworthy, especially compared to the number of questions on the employment prediction’s societal and political effects observed in other subcategories.
Much of participants’ information needs pertained to how the system would be used in practice (89 questions in total by over 20 participants). We differentiate questions on the system’s usage into everyday usage as a tool, usage as a sociotechnical component in human interaction, and usage embedded in a larger organizational structure. We include questions on misuse, costs, and examples, as they also address the system’s concrete application.
Usage as a tool: How is the system operated? Questions on practical details of usage were almost exclusively gathered in the wristband use case (30 of 31 total), with only one question emerging in the employment prediction use case. Questions typically concerned how the wristband would be used by family members, including waterproofing, charging, operation, and similar. Questions like these would likely be covered by a manual accompanying the wristband.
Usage as a sociotechnical component: How will the system impact interpersonal relations? Questions addressing how the systems would impact interpersonal work were gathered mostly from the employment prediction use case (13 of 14 total). Participants’ inquiries focused on how the system would affect interaction between consultant and job-seeker and covered both the consultant’s and job-seeker’s personal experience as well as their relationship:
Usage as part of an organization: How is the system integrated into existing structures? Questions that focused on the system’s interaction with existing organizational structures were present in both use cases, with a majority in the wristband use case (19 of 28 total). Questions on the institutional embedding covered a wide range of topics, including data processing, data access, information flow, contestability, dependence, and control: ”What are the options for correction? Is there a system loop that handles this?” (A7).
Examples: How would the system handle [this case]? Questions in this subcategory aimed to examine how the system would behave given a specific example case and emerged in low count in both use cases (6 total). Half of these were formulated as: ”How would the algorithm handle my case?” (A2).
Costs: What are the costs? Questions about costs pertained to both personal costs (”What would it cost me?” (B5)) as well as organizational costs (”If the state pays, what is the cost? Is the money well invested?” (A4)). This subcategory appeared more often in the wristband use case (6 of 8 total).
@@
Features: What features does the system consider and why? Questions on the system’s input features emerged almost exclusively in the employment prediction use case (28 of 31 total) and covered a range of different inquiry forms. Most questions on features emerged in the first inquiry phase but mirrored questions covered in the XAI Question Bank (Liao et al., 2020), e.g., Which features are considered?, How is a certain feature weighed?, and Why is this feature included? This points to shared information needs between ADM novices and experts, as questions on features in the question bank were sourced from stakeholders with technical knowledge. However, the use case appears to impact strongly whether features are inquired at all.
Functionality: What is the scope of the system? Can it do…? Questions on the system’s functionality emerged almost exclusively in the wristband use case (30 of 34 total) and in the first inquiry phase. The ’functionality’ subcategory differs from the ’usage as a tool’ subcategory in the inquiry’s focus: Questions on functionality inquired about the general existence or workings of specific functions (such as sleep monitoring or heart frequency measurements), whereas questions on the system’s usage as a tool inquired how the system would be operated in practice.
@@
Performance. Questions in this subcategory belonged to one of three categories which are covered by the XAI Question Bank (Liao et al., 2020): errors (What kinds of mistakes is the system likely to make?), limitations (What are the limitations of the system?), and reliability (How accurate/precise/reliable are the predictions?). While questions on reliability only emerged after reference to the question bank, questions on limitations and errors emerged organically and frequently even in the first inquiry phase: ”I mean, the question is, if the system really does make mistakes, how quickly can you give the feedback: ’Oops, that wasn’t a fall, everything’s okay.’ Is that possible?” (B4).
@@
Reflexive inquiry: What does [a machine learning concept] mean? This subcategory captures inquiries about the meaning of questions included in the XAI Question Bank (Liao et al., 2020), which emerged in both use cases (14 total). Often, these took the form of What does [a machine learning terminology] mean?, prompted by terms such as algorithm, biases, and predictions. However, a lack of understanding in vocabulary often correlated with a lack of understanding of the corresponding concept: ”’What kind of data was the system trained on?’ I don’t understand the ’trained’ part. That’s just an input, isn’t it?” (A3). We thus adapt the question slightly for this subcategory to ask about machine learning concepts.
4.2. RQ2-Understanding: How do question-driven explanations interact with participant understanding?
@@
This section provides an analysis of participant understanding: Section 8 describes the effects of inquiry and verbal explanations on self-reported understanding and decision confidence, and Section 4.2.2 describes how participant understanding interacted with question-driven explanations with a focus on explanatory stances and challenges in understanding.
4.2.1. How inquiry impacts understanding and confidence.
@@
Participants reported their perceived understanding and decision confidence on 5-point Likert scales before and after the inquiry phases.888One participant decided to not vote on the use case and did not answer the second query out of concern about how the study results would be used.. We describe three main aspects of these self-reports: the overall level of understanding, changes in understanding, and interactions between understanding and confidence.
| Employment prediction use case | A6 | A11 | A4 | A5 | A8 | A1 | A3 | A10 | A2 | A12 | A9 | A7 |
| Self-reported understanding | ++ | + | = | = | = | = | = | = | - | - | - - - | n/a |
| Confidence | + | = | + | - - | = | + | = | = | = | = | = | n/a |
| Health wristband use case | B8 | B10 | B2 | B4 | B5 | B6 | B7 | B9 | B12 | B1 | B3 | B11 |
| Self-reported understanding | + | + | = | = | = | = | = | = | = | - | - | - |
| Confidence | - | + | + | = | = | = | ++ | = | - | ++ | + | + |
Self-reported understanding was high overall, but ’high understanding’ differs between participants. 3 of 24 participants indicated ”3” or lower in their first self-reported understanding (3 = indifference between understanding and no understanding, 5 = no understanding at all) and only 1 participant indicated ”3” in their second report. However, when participants explained why they perceived their understanding to be high, the level of detail in their accounts differed noticeably. For example, domain expert A2 explained: ”The system should aid in deciding which measure to offer to a job-seeker. […] The question is how the algorithm is created, but the basic classification, I think, makes sense. (A2).
In contrast, participants with less domain expertise tended to provide less detailed explanations. B5, for example, referred to their reception of AI in the media as a reason for their high self-reported understanding: ”There is a lot of talk about artificial intelligence in the media. Mostly at night. I watch it when I can’t sleep.” (B5). Other participants spoke about the general employment market (A5), consequences of digitization (A9), and comparable systems (e.g., B8, A12). This illustrates a discrepancy in what people see as ”high” understanding, possibly resulting from the difficulty of assessing one’s understanding (Keil, 2006) (Section 5).
@@
Inquiry phases mostly did not influence self-reported understanding. After proceeding through the inquiry phases and the decision, most participants reported their understanding to have remained the same, 4 participants reported increased understanding and 6 reported declines (as seen in Table 3). As expected, increased self-reported understanding was explained through information gained in the inquiry phases. In contrast, decreases in understanding seemed to have various reasons, e.g., for A2 the decrease in understanding was accompanied by heightened scrutiny:
@@
Confidence increased for 9 participants but seems not to depend on understanding. In contrast to understanding, inquiry led to more increases than decreases in perceived decision confidence (as seen in Table 3). A final increase in confidence for several participants seemed to depend on how the system would work in practice: ”By and large, I know how the system works. Otherwise, to be sure, I would have to see how it actually works in practice.” (A4). Participants however offered little explanation as to why understanding did not increase together with their confidence.
@@
4.2.2. How question-driven explanations interacted with participant understanding.
@@
The cognitive sciences outline a number of processes governing how people seek and process information (described in Section 2). In this section, we analyze two main aspects of participant understanding: the role of explanatory stances and the challenges that prevented participants from effectively filling their gaps in understanding.
Explanatory stances guide how information is acquired. To analyze if participants preferred a specific way of seeking information, we apply the concept of the three explanatory stances (mechanical, design, and intentional) (Keil, 2006; Lombrozo and Wilkenfeld, 2019; Dennett, 1998).999A more detailed introduction is provided in section 2. Which stance is chosen guides how information is inquired and understood, e.g., inquiring information from a mechanical stance would prompt explanations about technical details. Examining which stance is applied in inquiry thus indicates if participants often used a particular stance for a particular topic. In the following, we describe which stances were applied most often in the question categories outlined in Section 4.1.
@@
Questions about data and system specifications often applied the mechanical stance. The mechanical stance focuses on parts, processes, and mechanisms and does not consider goals or intentions (Lombrozo and Wilkenfeld, 2019). Participants often applied the mechanical stance when asking about, e.g., features, predictions, and output. This likely resulted from reference to the XAI Question Bank (Liao et al., 2020), which mostly contains questions that elicit mechanistic and functional explanations, such as: ”What is the sample size of the training data?” (A4), ”How does the algorithm actually work?” (A3), and ”How often does the system make mistakes?” (B1). Small differences in question phrasing, however, can lead to different explanations. Compare: ”What is the sample size of the training data?” (A4) and ”The sample size of the data, I assume, is large enough?” (A5). While the first one is taken verbatim from the XAI Question Bank (Liao et al., 2020), the second raises the point: ”Large enough for what?”, which rather applies the design stance in asking for functional information.
Questions about the system’s context and usage often applied the design stance. The design stance focuses on how functions and practical goals are realized in terms of design (Keil, 2006). Participants tended to apply this stance to inquire about, e.g., practical intention, consequences, and usage as a tool or in an organization. Typical design inquiries included: ”What’s the benefit of knowing whether the patient was at the refrigerator at 3:00 or in the bathroom at 2:00?” (B4), and ”What if someone is lying in bed and now has a heart attack? Can the device detect that, too?” (B5). Notably, questions on functionality almost exclusively emerged in the wristband use case.
Questions about discrepancies in values applied the intentional stance. The intentional stance focuses on underlying purpose and connects explanations to the beliefs and desires involved (Keil, 2006; Dennett, 1998). In both use cases, participants applied this stance when asking about the structural intentions of the deploying institutions: ”Would that mean that the employees of the job agency would need to perform better?” (A4). Participants also applied this stance when inquiring about their rights as affected stakeholders: ”If I’m analyzed, I would like to have the results of the analysis. Is there such a thing as a right of appeal?” (A8). Combined, this indicates that the intentional stance was applied when participants thought their own values would not be considered in the system, which seemed to occur more often in the employment prediction than the wristband use case.
@@
Challenges in acquiring information to develop understanding. Explanations bring the incompleteness of one’s understanding into focus and prompt specific reactions: filling the gap with further information, ”outsourcing” it to another mind, or mentally rendering it irrelevant (Keil, 2006). The explanation setting used in this study explicitly supported filling gaps with information by letting participants flexibly inquire about any topic. However, participants also met challenges in acquiring information and developing understanding. In the following, we describe three of these challenges:
@@
Blind spots inhibit further inquiry. In the closing interviews, multiple participants expressed that they did not understand all relevant aspects of the system but could not articulate which information was missing. Reasons given for this effect included a lack of mathematical knowledge (A6), too little preparation (A8), pressure of the study situation (B5, A12), and uncertainty about the system’s practical behavior (A1, A4, B1, B9): ”Experience has taught me that it’s always the practical application that raises issues.” (B1). We call understanding gaps that could not be closed ”blind spots”.
@@
Outsourcing understanding divides cognitive labor but risks over-dependence. Multiple participants asked about other people’s judgment of the system, e.g., the job center’s employees (A8), digital natives (A9), their families (B11), and the study examiner (A11). This ”outsourcing” (Keil, 2003) is a common process to divide the cognitive labor of understanding between trusted people. Critically, some participants stated intention to outsource their whole understanding: ”The first question is: Would you have said yes to this system?” (A9). Trust is thus a crucial part of stakeholders’ understanding when outsourcing is used, especially as it risks over-dependence on the entrusted person.
@@
Abandoning of understanding undermines the explanation effort. Some participants expressed that they would likely not care about the system even if it was used on them, resulting from a lack of interest or the belief that their voice would be irrelevant: ”I still think that it’s not my interest that is being represented here, but actually the interest of the job agency.” (A8). A pronounced disinterest in topics despite their relevance for oneself is known in political science as ”civic ignorance” (Lupia, 2016)101010The term describes that ignorance on public matters applies to almost everyone (Lupia, 2016).. Research has been conducted on ways to remedy this disinterest and will be briefly discussed in Section 5.
@@
Participants were asked for their perceptions of risks and benefits before inquiry and after making the decision (Section 3). Changes in their perceptions from the first to second report are depicted in Figure 3, showing strong changes in the employment prediction use case and less so in the health wristband use case. In the following, we provide a brief overview of participants’ perceptions in both use cases.
Perceptions changed strongly in employment prediction use case. Participants in the employment prediction use case showed diverse initial perceptions (as seen in Figure 3) and strong changes in their perceptions after the inquiry phases. Perceptions were further likely influenced by the participants’ familiarity (Kramer et al., 2018) with the AMS algorithm, which drew considerable public attention during its piloting phase in 2018 (Allhutter et al., 2020a). Frequently mentioned benefits included a potentially reduced workload, debureaucratization, and guidance for personnel consultants, while frequently mentioned risks included a lack of care for the individual, fear of discrimination, and overreliance on algorithmic decisions. A6, for example, reported a decrease in perceived societal risk and benefit, while also reporting an increase in perceived personal benefit, explaining:
My intuition is that an algorithm like this can make my work a lot easier because it gives me guidance. But if it’s used too much as a guideline, then it naturally restricts the options of the people I look after. (A6)
@@
In contrast, A12 reported a strong decrease in perceived risk and increased benefit in both dimensions, stating:
Honestly, I haven’t heard anything bad. It can only have positive benefits. […] I say it’s worth a try. And in any case, an assessment will be made based on the data that is available. (A12)
@@
In total, we observe mixed effects that question-driven explanations had on participants perceptions in the employment prediction use case. While the provided information certainly enabled a shift in perspective, the trajectory of that shift seems to differ heavily, even though the information provided did not change between participants except, of course, if different questions were asked.
Perceptions changed less in the health wristband use case. Participants in the health wristband use case reported less pronounced changes in perceptions and overall saw the use case rather clearly as one of low risk and high benefit. Often participants weighed the risk of surveillance against the benefit of the wearer’s safety, as B11 explains:
@@
For me I don’t see any disadvantages. But I could imagine that there are people who say: ”I don’t want to be monitored. Even if I fall down and nobody comes and I bleed to death, then I’ll just bleed to death.” (B11)
B12 assigned the system a higher societal risk but also a high benefit, concluding:

### v6

@@ 4. Results
This category captures questions about the data used by the system. These questions emerged in both use cases (43 questions total) and mostly occurred only after participants received the XAI Question Bank (Liao et al., 2020) as reference. The question bank already contains questions about data type (What kind of data was the system trained on?) and data source (What is the source of the training data?). In our study, multiple participants also asked about data privacy and misuse (11 questions by 8 participants). Therefore, we extend the data category with a subcategory on data safety (Are the data correct / representative / safe?).
@@
Features: What features does the system consider and why? Questions on the system’s input features emerged almost exclusively in the employment prediction use case (28 of 31 total). Most questions on features emerged before participants were given the XAI Question Bank (Liao et al., 2020) but mirrored questions covered in the bank, e.g., Which features are considered?, How is a certain feature weighed?, and Why is this feature included? This points to shared information needs between AI novices and experts, as questions on features in the question bank were sourced from stakeholders with technical knowledge. However, the use case appears to have a strong impact on whether features are inquired at all.
Functionality: What is the scope of the system? Can it do…? Questions on the system’s functionality emerged almost exclusively in the wristband use case (30 of 34 total). The ’functionality’ subcategory is close to the ’usage as a tool’ subcategory but differs in its focus: Questions in this subcategory asked about functional scope in general (“How does the wristband work?” (B5)), whereas questions in the “usage as a tool” subcategory“ instead asked how the system would be used specifically and practically (”Does that mean you can also shower with it?“ (B4)).
@@
Errors, limitations, reliability Questions about errors, limitations, and reliability (21, 23, and 7; 51 total) emerged in both use cases equally and covered topics such as error frequency and susceptibility, handling incomplete data, constraints of usage scenarios, and precision. While questions on reliability only emerged after reference to the bank, questions on limitations and errors emerged organically even in the first inquiry phase: ”I mean, the question is, if the system really does make mistakes, how quickly can you give the feedback: ’Oops, that wasn’t a fall, everything’s okay.’ Is that possible?“ (B4).
@@
Reflexive inquiry: What does [a machine learning concept] mean? This subcategory captures questions about the meaning of certain notions in machine learning, which emerged in both use cases (14 total) after participants read through the XAI Question Bank (Liao et al., 2020). Questions were often prompted by terms such as algorithm, biases, and predictions. However, a lack of understanding in vocabulary often correlated with a lack of understanding of the corresponding concept: ”’What kind of data was the system trained on?’ I don’t understand the ’trained’ part. That’s just an input, isn’t it?“ (A3). We thus adapt the original bank’s question slightly such that it captures questions about machine learning concepts instead of terminologies.
4.2. RQ2-Understanding: How do question-driven explanations interact with participants’ understanding?
@@
This section describes how the explanations given in response to participants’ questions interacted with their understanding and decision confidence. First, we focus on the effects of question-driven explanations on self-reported understanding and decision confidence (Section 9). Second, we describe how participants’ questions can be analyzed using the framework of ”
4.2.1. How question-driven explanations impact understanding and confidence.
@@
Participants reported their perceived understanding and decision confidence on 5-point Likert scales before and after the inquiry phases, i.e., before and after receiving explanations (overview depicted in Table 3).999One participant decided not to vote on the use case and did not answer the second query out of concern about how the study results would be used. We describe three main aspects of these self-reports: participants’ overall perceived understanding, how their understanding changed throughout the study, and how their understanding interacted with their decision confidence.
Table 3. Change in self-reported understanding and decision confidence. Participants indicated their perceived understanding of the system and confidence in their decision on a 5-point Likert scale before and after receiving question-driven explanations in the inquiry phases (Section 3.2). This table summarizes the change between their two self-reports. Notably, changes in perceived understanding did not necessarily lead to increases and decreases in confidence. Changes between reports are encoded with symbols and colors, + (increase), - ( decrease), and = ( no change). Columns indicate participant ID and use case.
Employment prediction use case
A6
A11
A4
A1
A8
A3
A10
A5
A2
A12
A9
A7
Self-reported understanding
n/a
Confidence
n/a
Health wristband use case
B10
B8
B7
B2
B4
B5
B6
B9
B12
B1
B3
B11
Self-reported understanding
Confidence
Self-reported understanding was high overall, but participants define ’high understanding’ differently. In the first self-reported understanding, before receiving explanations, only 4 of 24 participants indicated that they did not understand the ADM system well. In the second self-report, after receiving explanations, there even was only 1 participant who indicated that they did not understand the system and who perceived a substantial decrease in their understanding (”
Question-driven explanations mainly did not influence self-
@@
reported understanding. After proceeding through the inquiry phases, most participants reported their understanding to have remained the same; 4 participants reported increased understanding, and 6 reported declines (as seen in Table 3). As expected, increased self-reported understanding was explained through information gained in the inquiry phases. In contrast, decreases in understanding seemed to have various reasons, e.g., for A2, the decrease in understanding was accompanied by heightened scrutiny:
@@
Confidence increased for 9 participants but seems not to depend on understanding. Compared to understanding, question-driven explanations increased participants’ decision confidence more often (as seen in Table 3). Several participants noted that it would increase their confidence further if they saw the system in practice: “By and large, I know how the system works. Otherwise, to be sure, I would have to see how it actually works in practice.” (A4). Participants, however, offered little explanation as to why understanding did not increase together with their confidence.
@@
4.2.2. How question-driven explanations interacted with explanatory stances and which challenges arose.
@@
The cognitive sciences outline several processes governing how people seek and process information (described in Section 2). In the following, we analyze two main aspects of participant understanding: the role of explanatory stances and the challenges that prevented participants from acquiring a “working understanding” (Keil, 2006).
Explanatory stances offer a way of matching explanations to participants’ information needs. Explanatory stances are a way of describing how people acquire information to develop a predictive strategy of a system’s (or person’s) behavior (Dennett, 1998). To this end, people can take different strategies, which might be focused on a system’s physical attributes, on its design and functions, or, if those are inaccessible, on the “beliefs [an] agent ought to have given its place in the world and its context” (Dennett, 1998), which define its intentions and actions.
To analyze if participants preferred a specific strategy of seeking information, we apply the concept of the three explanatory stances (mechanical, design, and intentional) to their questions during the inquiry phases.101010An introduction is provided in section 2. We conduct this analysis by assigning participants’ questions the likeliest of the three stances. This assignment is an interpretation of their chosen stance derived from the question’s phrasing and the conversation’s context. Consequently, several questions might be assigned to more than one stance, depending on the interpretation of the information needs (discussed in Section 5.2.2). In the following, we describe our analysis of which stances participants applied for their questions about the ADM systems:
@@
Questions about data and system specifications often applied the mechanical stance. The mechanical stance focuses on parts, processes, and mechanisms and does not consider design or intentions (Lombrozo and Wilkenfeld, 2019). Participants tended to apply this approach to learn about the technical details of the system, such as the data basis, features, predictions, and output. This direction of inquiry often resulted from reference to the XAI Question Bank (Liao et al., 2020), which contains many questions about technical details, including “What is the sample size of the training data?” (A4), “What are the features with the evaluation criteria?” (A7), and “How sensitive is it?” (B8). However, small differences in question phrasing can lead to different explanations. Compare: “What is the sample size of the training data?” (A4) and “The sample size of the data, I assume, is large enough?” (A5). While the first one is taken verbatim from the XAI Question Bank (Liao et al., 2020) and can be said to focus on a “part” of the system (the training data), the second raises the question of what the sample size should be large enough for, which rather applies the design stance by asking about functional information.
Questions about the system’s usage and context applied the design stance. The design stance focuses on how functions and goals are realized in terms of design (Keil, 2006). Therefore, we assign participants’ questions about the system’s design’s practical and conceptual aspects to this stance. Examples of practical design aspects include questions about usage details and operation instructions (“Do you wear it day and night?” (B1)). In contrast, examples of conceptual design aspects include questions on the motivation behind certain functionalities (“What’s the benefit of knowing whether the patient was at the refrigerator at 3:00?” (B4)). Further, we assign questions about context information to this stance when they acquire information about, e.g., the characteristics of the target group (“This thing is only supposed to be distributed to care home residents, right?” (B12)) or control mechanisms (“What options for correction are there?” (A7)).
Questions about context and values applied the intentional stance. The intentional stance describes a system’s behavior based on beliefs and motives embedded in it, which are assumed (in the absence of mechanical and design information) to drive its actions (Dennett, 1998). We assigned participants’ questions to this stance when they addressed topics such as the intentions of the deploying institutions (“Would that mean that the job counselors would need to perform better?” (A4)) and the rights of affected persons (“If I’m analyzed, I would like to have the results. Is there such a thing as a right of appeal?” (A8)). Further, the intentional stance also tended to be used when participants suspected that their values would not be considered in the system (“When this comes, who guarantees that it is not forced on you?” (B6)). Notably, many questions in the intentional stance focused on the beliefs and values of the deploying institution, which connects to literature emphasizing the importance of people’s perceptions toward the institution (Cavaliere and Romeo, 2022). This showed that participants regarded the job agency critically and is reflected in participants’ perceptions of the systems’ risks and benefits (Section 4.3).
@@
Challenges in acquiring information to develop understanding. Explanations bring the incompleteness of one’s understanding into focus and prompt specific reactions: filling the gap with further details, “outsourcing” it to another mind, or mentally rendering it irrelevant (Keil, 2006). The explanation setting used in this study explicitly supported filling gaps with information by letting participants flexibly inquire about any topic. However, participants also met challenges in acquiring information and developing understanding. In the following, we describe three of these challenges:
@@
Blind spots inhibit further inquiry. In the closing interviews, multiple participants expressed that they did not understand all relevant aspects of the system but could not articulate which information was missing. Reasons given for this effect included a lack of mathematical knowledge (A6), too little preparation (A8), pressure of the study situation (B5, A12), and uncertainty about the system’s practical behavior (A1, A4, B1, B9): “Experience has taught me that it’s always the practical application that raises issues.” (B1). We call understanding gaps that could not be closed “blind spots”.
@@
Outsourcing understanding divides cognitive labor but risks over-dependence. Multiple participants asked about other people’s judgment of the system, e.g., the job center’s employees (A8), digital natives (A9), their families (B11), and the study examiner (A11). This “outsourcing” (Keil, 2003) is a common process to divide the cognitive labor of understanding between trusted people. Critically, some participants stated intention to outsource their full understanding: “The first question is: Would you have said yes to this system?” (A9). Trust is thus a crucial part of stakeholders’ understanding when outsourcing is used, especially as it risks over-dependence on the entrusted person.
@@
Abandoning of understanding undermines the explanation effort. Some participants expressed that they would likely not care about the system even if it was used on them, resulting from a lack of interest or the belief that their voice would be irrelevant: “I still think that it’s not my interest that is being represented here, but actually the interest of the job agency.” (A8). A pronounced disinterest in topics despite their relevance for oneself is known in political science as “civic ignorance” (Lupia, 2016)111111The term describes that ignorance on public matters applies to almost everyone (Lupia, 2016).. Research has been conducted on ways to remedy this disinterest and will be briefly discussed in Section 5.
@@
Participants were asked how they perceived the risks and benefits of the ADM system before receiving explanations and after making the deployment decision. The two-dimensional grids used to elicit these perceptions divide the space into four quadrants, corresponding to four combinations of risk and benefits (low-risk low-benefits, low-risk high-benefits, high-risk high-benefits, and high-risk low-benefits; Section 3.2.4). Changes in their perceptions between the first and second report are visualized in Figure 4, showing solid changes in the employment prediction use case and lesser changes in the health wristband use case. In the following, we describe participants’ perceptions in both use cases and outline potential reasons for the changes.
Perceptions changed enormously in employment prediction use case. Participants in the employment prediction use case showed very different initial perceptions of the system, ranging over all combinations of risks and benefits and showing little focus on specific attitudes (as seen in Figure 4). Further, participants enormously changed their perceptions of the ADM systems after receiving explanations, often transitioning into other perception quadrants due to newfound information.
A6: What does “regional labor market” mean? That a large number of jobs are registered as vacant?
Examiner: Yes, the city was divided into regional labor market types. It was intended to show how many jobs are available, but it is also a placeholder variable for the place of residence.
A6: I find that quite discriminatory, I have to say.
Frequently mentioned benefits included a potentially reduced workload, reduced administrative overhead, and guidance for personnel counselors, while frequently mentioned risks included a lack of care for the individual, fear of discrimination, and overreliance on algorithmic decisions. A11, for example, reported an increase in perceived societal benefit, explaining:
I believe that the state could save money in the broadest sense. And this money could be used to plug other holes. […] It would only be a risk because soft skills and the emotional part are not included. That’s the only risk I can see. (A11)
@@
In contrast, A2 reported a slight decrease in perceived risk and varying changes in benefit while also voting “no”, stating:
What I’m missing here is that no consideration is given to whether the person looking for work can continue in their core business or needs a new career path. […] I miss the jobseeker’s personal perspective. The hard facts are the smaller issue for me. (A2)
@@
We observed that question-driven explanations had mixed effects on participants’ perceptions in the employment prediction use case. While the provided information indeed enabled a shift in perspective, the trajectory of that shift seems to differ heavily, even though the information provided did not change between participants.
Perceptions changed less in the health wristband use case. Participants in the health wristband use case saw the system clearly as having low risk and high benefit. Changes in perceptions were minor and never resulted in transitions between perception quadrants, showing that participants perceived the system positively before and after receiving explanations.
Through the question and answer sessions, my idea of the wristband has become more concrete. I realized that the thing can do other things than I had imagined. But with risks and benefits, I’m sticking with what I had before, that hasn’t changed. (B7)
For some participants, the wristband presented a trade-off between the values of safety and privacy, which they weighed against each other. Despite little overall change in perceptions, this highlights that participants considered the risk of surveillance and the benefit of the wearer’s safety and came to different conclusions. B11, for example, personally did not see any issues, while B4 specifically mentioned the feeling of being surveyed:
@@
For me, I don’t see any disadvantages. But I could imagine that there are people who say: “I don’t want to be monitored. Even if I fall down and nobody comes and I bleed to death, then I’ll just bleed to death.” (B11)
I would rate the risk as high because I believe it’s going in the direction of surveillance. (B4)
In summary, participants thus weighed the given information and decided based on it, which suggests that the explanations fulfilled their purpose in revealing potential value conflicts in the system’s deployment. Making this conflict explicit, B12 stated: